the perimeter of a rectangle is 86m. The length is 7m more than the width. Find the dimensions.
Answer by mananth(16946)

the perimeter of a rectangle is 86m. The length is 7m more than the width. Find the dimensions.
Let width be x
Length = x+7
Perimeter= 2*(l+w)
=2*(x+7+x)
=4x+14= 86
4x= 86-14
4x=72
X=18
Length =25 meters
mananth@hotmail.com